A Bibliography of Publications on Fractal Image Compression

Brendt Wohlberg University of Cape Town Electrical Engineering

Cape Town South Africa

Tel: ?n/a? FAX: ?n/a?

E-mail: brendt@dip.ee.uct.ac.za (Internet)

29 January 2020 Version 1.19

Abstract

This bibliography records publications on fractal image compression.

Title word cross-reference

7th [Ano94].

’88 [IEE88].

Accelerating [Sau95a]. acceleration [SH96]. Accurate [SLN95]. Acoustics [IEE93]. acquisition [EW96]. Adaptive [BV93, BV94, Dav95a, DBC93a, DBC93b, DC94b, DC94a, HMS93, LØ95, LLP96, Reu94a, SLTC94]. Address [VI93]. Advances [Bea90]. Affine [CD93, Hol91, DV93]. Affinity [Woh96]. Algorithm [BJ95, GK88a, GK88b, Kom95a, LV95, Sho89, CK96b, GAH93, TW94, Woo94]. Algorithms [FV94b, Lep93, MD95, VR94, RBM96]. Analysis [Ano95, CASH92, Sig96]. Application [BMK93, Bog95a, Fis95, Bon96]. Applications [BMK95, Bog95b, LUW95, ØBL^{+}94, FL92]. Approach [CFMV92, ØLR91, HB93, PH91]. Approximation [CZ97, FV94a, FV94b, KP94, KP96, MD92a]. April [IEE93, SC93, SC95]. Archetype [BJ95]. ASI [Ano95]. Assumption [Woh96]. Attractor [DC94b, DC94a, Lep93, LØ95, Nov93a, Nov93c, Nov93b]. Automata [CK93].

Barnsley [Gra92]. Based [Bea90, CASH92, Dav95a, Dav95b, DACB96, Det92, FGHS94, GAH94, GK88a, GK88b, Jac92, JR90, KNH94, KP94, KP96, Pen84, Pen94, Reu94a, TD95, ZY91, BM92, Che95, EW96, GAH93, Jac90a, KPY96, LNF93, OCK95, PH91, RFT94, Reu93, Reu94b, SLTC94, Sig95, TW94, TTMB96, WY96, WHK93]. Basis [FV94a, Vin95]. Behind [DK89]. Better [BS88]. Between [Sim95a]. bit [Slo94]. Block [Bea90, Jac90b, LB94, MD92b, MW94b, RC94, RC95, WdJ95, CK96b, GAH93, KPY96, Liu95a]. Block-Coding [Jac90b, GAH93]. Blocks [MD92a]. Bounds [Hür95a]. Breaking [Sau94]. Buffered [WNM94].

California [LUW95]. Center [IEE93]. centroid [Bon95]. Chaos [BJM^{+}88, DK89, PJS92a]. Chaotic [BD86]. Characteristics [KCC87]. Class [Mon93a, ØL95]. Classification [BJ95]. Clustering [LØ95, Ham95]. Codebook [Ham95, Kom95b, LØ95, HS93, Sig96]. Coders [ØL95]. Coding [BMK93, BMK95, BV93, BVN93, BV94, BSVN94, Bar95, BV95, Bea90, BDK92, Bog94a, Bog94b, Bog95a, Bog95b, DC94b, DC94a, DSC95, DV94, Dud94, GAH94, GK88a, GK88b, HMS93, HM94, HS94, Hür95a, Jac90b, Jac92, Jac93, JR90, KNH94, KP94, KP96, LB94, Lin94, MD92b, Nov93a, Nov93c, Nov93b, ØLR91, ØBL^{+}94, Øie94, PDY97, Reu93, Reu94a, Reu94b, RC94, RC95, WM95, BBBCP96, BM95, BM92, Bon96, CK96b, GAH93, GAH96, GIG95, Hür93, HB93, HS93, HMS94, Hür95b, Jac90a, Liu95a, RBM96, SLTC94, Sig95, Sim95b, Slo94, TTMB96, WY96, WdJ96, WHK93, ZZZ^{+}96]. Collage [DV94, ØBL^{+}94]. Color [GK88a, GK88b, HMS94, ZP95]. Communications [Bog95b, IEE88, MN94]. Comparative [DBC93a, DBC93b]. Comparison [EW96, FSR94]. Comparisons [Lep93]. compensation [BM95]. Complexity [Mon93b, Reu94b, Sau94, SH94]. Compress [BS88]. Compression [Ans93, BJ95, CASH92, CHF94, CHF95, CK93, Dav95a, Dav95b, DBC93a, DBC93b, DACB96, Det92, DV94, Fis92a, Fis92b, FJB92, Fis95, FGHS94, GK88a, GK88b, Hol91, Kom95a, Lep93, LV95, LY94a, LY94b, MW94b, Sau94, SH94, Sau95a, TD95, VR94, vdW95, vdW96, WdJ94, WdJ95, Woh96, ZY91, BE95a, BE95b, Bon95, CK96a, Che95, Dav96, DV93, EW96, FL92, Ham95, HMS96a, HMS96b, Har94, KKL95, Kom95b, KPY96, LNF93, LV96, Liu94, LLP96, MW94a, MSK^{+}95, OCK95, PH91, RFT94, Sau95b, SR96, Sau96a, SH96, Sau96b, Sig96, SC93, SC95, TW94, Woo94, WM94, ZP95, ZY94]. Computations [FV94b]. Computer [DK89, HCF95]. Computing [Sho89]. Conference [IEE88, IEE93, IEE95, Ano94, SC93, SC95]. Constrained [Bar95]. Construction [Dud94, HM90]. Contraction [BDK92]. Contractive [Jac92, ØLR91, Jac90a]. Contractivity [Hür93]. Convention [IEE93]. Convergence [HH94, HS94, Kom95c, ØL95]. convolution [SH96]. cosine [FSR94, ZY94]. coupling [SLTC94].

Data [GK88a, GK88b, SC93, SC95, KPY96, Woo94]. DC [IEE95]. DCC’93 [SC93]. DCC’95 [SC95]. DCT [FSR94, WdJ95]. Decoder [ØL95, KKL95]. Decoding [BMK93]. Decomposition [RC94, WHK93]. Delaunay [DC94a, DC94b, DACB96]. Description [CD93, Pen84, Pen94]. Detection [CASH92]. Deterministic [MD95]. Diego [LUW95]. Difference [Bog94a]. Digital [Hol91, IEE88, Jac90b, LB94]. Dimension [KCC87]. Dimensional [BV95, Hol91, VI93, Sau95a, ZCT94]. Discrete [KMK95, Lun95, MH91, MH92, FSR94, ZY94, ZCT94]. Discussion [Fis92a]. Distance [Sho89]. Distortion [MW94b, GAH96, WM94, ZP95]. Domain [BV93, BV94, WdJ95, Sau96b]. Dynamics [BD86].

Edge [CASH92]. Edinburgh [Ano94]. Efficiently [Sho89]. Encoded [Kom95c]. Encoding [Ano95, FRS94, Gra92, LØ95, PJS92b, WdJ94, COK96]. enhanced [HMS96b]. Enhancing [BE95a, HMS96a]. Entropy [Bar95]. Equation [Bog94a, Liu95b]. equations [Liu93b, Liu95a]. ergodic [Elt87]. European [Ano94]. EUSIPCO’94 [Ano94]. evaluation [Hür95b]. Evolutionary [SR96]. exact [Liu95a]. Explicit [Sim95a]. Expression [CD93]. Extensions [Liu93a].

Fast [BMK93, COK96, Dud94, HS93, Kom95a, KPY96, Lep93, LØ95, LV95, ØL95, WdJ95, RBM96, SH96, TW94]. faster [Bon95]. features [GIG95]. Fifth [SC95]. Finite [CK93, CK96b]. finite-state [CK96b]. first [WdJ96]. first-order [WdJ96]. Form [Det92]. Fractal [AT89, ATD92, Ano95, Ans93, BMK93, BMK95, Bar88, Bar93, BH93, BV93, BVN93, BV94, BSVN94, Bar95, BV95, Bea90, BDK92, BBBCP96, Bog94a, Bog94b, Bog95a, CK96b, Che95, CZ97, CASH92, CHF94, CHF95, Dav95a, Dav95b, DBC93a, DBC93b, DSC95, DACB96, Det92, DV94, Dud94, Fis92a, Fis92b, FL92, FJB92, FRS94, Fis95, FGHS94, GAH94, GAH96, GK88a, GK88b, GIG95, Gra92, HM90, Har94, Hol91, HMS93, HB93, HH94, HMS94, HM94, HS94, Hür95a, Jac90a, Jac90b, Jac92, Jac93, KCC87, KP94, KP96, Kom95a, Kom95c, KMK95, LB94, LNF93, Lin94, LV95, Liu93b, Lun95, MH91, MD92a, MD92b, Mon93a, Mon93b, Mon93c, MW94a, MW94b, MN94, MSK^{+}95, OCK95, ØBL^{+}94, Øie94, ØL95, PS88, Pen84, Pen94, PDY97, RFT94, Reu93, Sau94, SH94, Sau95a, Sau95b]. Fractal [SLTC94, Sim95a, SLN95, TW94, TD95, VR94, VI93, vdW95, vdW96, WNM94, WdJ94, WdJ95, WdJ96, Woh96, WHK93, WM95, ZY91, ZP95, ZCT94, BE95a, BE95b, COK96, CK96a, Dav96, DV93, EW96, FSR94, FV95, GAH93, Ham95, HMS96a, HMS96b, Hür93, HS93, Hür95b, KKL95, Kom95b, KPY96, LV96, Liu93a, Liu95a, Liu95b, LLP96, PH91, RBM96, SR96, Sau96a, SH96, Sau96b, Sig96, Slo94, TTMB96, WY96, Woo94, WM94, ZZZ^{+}96]. Fractal-Based [Det92, GAH94, Pen84, Pen94, LNF93, OCK95, RFT94, TW94, WHK93, EW96, GAH93, PH91]. fractal/wavelet [CK96a]. Fractals [BD86, CFMV92, CL93, DK89, JR90, MD95, PJS92a, Zor88, HCF95, ZY94]. Frame [Bog94b]. Framework [Lun95]. Franklin [IEE88]. Frequency [BV93, BV94]. Frontiers [PJS92a]. Function [Det92, FV94a, FV94b, LY94a, LY94b, MH92, PD95, Vin93, Har94, WY96]. Function/Image [FV94a, FV94b]. Functions [Hol91, Reu94a, Reu94b]. futility [Sau96a]. Fuzzy [CFMV92].

Gauss [WdJ96]. General [Woo94].

Generalized

[GAH94, Lin94, Mon93b, FV95]. Genetic [RBM96, VR94]. Geometrical [Sig95, Sig96]. Geometry [GK88a, GK88b, ZY91]. Graphics [DK89].

Harnessing [BJM^{+}88]. Hashing [HCF95]. Hausdorff [Sho89]. Heuristic [TD95]. Hidden [MH91]. Hidden-Variable [MH91]. Hierarchical [BMK93, BMK95, Dud94, HS93]. Hotel [IEE88]. Human [JR90, EW96, MSK^{+}95]. Hybrid [Mon93c, WM95, ZY91, ZZZ^{+}96, RFT94].

ICIP’95 [IEE95]. IFS [Sig95, Vin95]. II [FV94b]. III [LUW95]. Image [Ano95, Ans93, BMK93, BMK95, BJM^{+}88, BV93, BVN93, BV94, BSVN94, Bar95, BDK92, Bog94a, Bog95a, Bog95b, BJ95, CASH92, CK93, CD93, Dav95a, Dav95b, DBC93a, DBC93b, DC94b, DC94a, DSC95, DACB96, Det92, Dud94, Fis92a, Fis92b, FJB92, Fis95, FV94a, FV94b, FGHS94, GAH94, Hol91, IEE95, Jac92, Jac93, JR90, KP94, KP96, Kom95a, LUW95, Lep93, LØ95, Lin94, LV95, Liu95a, LY94a, LY94b, MD92a, ØLR91, ØBL^{+}94, Øie94, ØL95, PDY97, Reu94a, Reu94b, RC94, RC95, Sau94, SH94, Sau95a, Sho89, TD95, VR94, vdW95, vdW96, WdJ94, WdJ95, Woh96, WM95, ZY91, ZY94, BE95a, BE95b, BM95, BM92, Bon95, Bon96, COK96, CK96a, Che95, CL93, Dav96, EW96, FL92, GAH93, GAH96, GIG95, Ham95, HMS96a, HMS96b, Har94, Hür93, Jac90a, KKL95]. image [Kom95b, KPY96, LNF93, LV96, Liu94, LLP96, MW94a, MSK^{+}95, OCK95, PH91, RBM96, Sau95b, SR96, Sau96a, SH96, Sau96b, SLTC94, Sig95, Sig96, Sim95b, Slo94, TW94, TTMB96, WY96, WHK93, WM94, ZP95, ZZZ^{+}96]. Images [BS88, GK88a, GK88b, Gra92, Jac90b, Kom95c, KMK95, MD92b, Nov93a, Nov93c, Nov93b, PS88, PJS92b, SLN95, VR94, Bar88, Bar93, BH93, BBBCP96, HS93, HMS94, Woo94]. implementation [ZCT94]. Implicit [Dav96]. Improved [BVN93, ØBL^{+}94]. Inner [ØLR91]. Inter [Bog94b]. Inter/Intra [Bog94b]. Inter/Intra-Frame [Bog94b]. Interband [Sim95b]. International [IEE88, IEE93, IEE95]. Interpolation [Hol91, MH91]. Interpretation [BMK93, BMK95, Sig95]. Intra-Frame [Bog94b]. invariant [GIG95]. Inverse [AT89, ATD92, CFMV92, FV94a, FV94b, HM90, Har94, HCF95]. isometries [Sau96a]. Issue [Reu94b]. Issues [Mon93b]. Iterated [BJ95, CFMV92, FJB92, FV94a, FV94b, Jac92, KNH94, LY94a, LY94b, MH92, PD95, Reu93, Reu94a, Reu94b, Vin93, Elt87, Har94, Jac90a, WY96]. Iteration [BM92]. Iterative [Bog95b, DV94, KKL95].

JPEG [EW96]. July [Ano95, LUW95].

June [IEE88]. Just [Zor88].

Kohonen [BM92].

Lean [Sau96b]. Length [GK88a, GK88b]. limitations [DV93]. Line [AT89]. linear [HCF95]. Link [Sim95a]. Local [ATD92, Sim95a]. Lossless [SH96, LV96]. Low [Slo94, HB93]. Low-bit-rate [Slo94].

magnification [GIG95]. Maps

[VI93, Elt87, Ham95]. March [SC93, SC95]. Markov [WdJ96]. mass [FL92]. Matching [WdJ95, GAH96, KPY96]. Mathematics [DK89]. measure [ZP95]. Measures [ATD92]. medical [BBBCP96]. Merging [vdW96]. Method [AT89, GAH94, HM90, ZY91, KKL95, KPY96, OCK95, Sim95b, ZZZ^{+}96]. Methods [SH94, vdW95, vdW96, COK96, FSR94, Liu94]. Minneapolis [IEE93]. Minnesota [IEE93]. Mixed [DSC95]. Model [MH92, PDY97, VI93, LLP96, ZCT94]. Modeling [Lun95]. Modelling [HM94, Vin93, Woh96, Bar88, Bar93, BH93]. models [Dav96, KPY96, WdJ96]. Moment [AT89, HM90]. Moments [ATD92]. Multi [KNH94, Sau95a]. Multi-dimensional [Sau95a]. Multi-resolution [KNH94]. Multiresolution [Bon96, CZ97, ØBL^{+}94, RC95, Sim95a, Bon95]. Multiscale [Bog94a, Bog94b].

NATO [Ano95]. Natural

[KCC87, Pen84, Pen94]. Nearest [Sau95a, Sau95b]. Neighbor [Sau95a, Sau95b]. network [BM92]. neural [BM92]. non [KKL95]. non-iterative [KKL95]. Nonlinear [PDY97, VI93]. Norway [Ano95]. Novel [Jac90b, KKL95]. observers [EW96]. October [IEE95]. One [VI93, ZCT94]. One-Dimensional [VI93, ZCT94]. operator [Bon95]. Operators [BDK92]. optimized [GAH96]. Optimum [WM95]. Order [MW94b, WdJ96]. organizing [Ham95]. Orthogonal [Vin95]. Other [CFMV92]. Overlapped [Reu94a].

PA [IEE88]. Parameter [Øie94]. Parameters [WM95]. Partial [WY96]. partially [DV93]. Partition [DSC95]. Partitioned [LY94a, LY94b]. Partitioning [Reu94a, Reu94b]. pattern [Bon95]. perception [Che95]. Perceptually [LV96]. Performance [Hür95a, WdJ96, WM94]. Personal [MN94]. Perspective [Woh96]. Philadelphia [IEE88]. Picture [Zor88]. Pictures [Bea90, HMS93]. PIFS [LLP96]. Plaza [IEE88]. Polynomial [BM95]. polyphase [WHK93]. pools [Sau96b]. Practical [PH91, SLN95]. Prediction [RC94, RC95, Sim95b]. Predictive [Lin94]. Pretty [Zor88]. Problem [ATD92, CFMV92, FV94a, FV94b, HS94, Har94, HCF95]. Problems [AT89, HM90]. Proceedings [Ano94, IEE95]. Processing [IEE93, IEE95, LUW95, Ano94]. Product [ØLR91]. pursuit [GAH96]. Pyramid [Bog95a, LV95, RC94].

Quadrilateral [DSC95]. Quadtree [LY94a, SLTC94]. quantised [BM95]. Quantization [Dav95a, DACB96, HMS96a, KP94, KP96, Øie94]. Quantized [Dav95b].

Rate [MW94b, WNM94, WM94, GAH96, HB93, Slo94]. Rate-distortion [WM94, GAH96]. Rate/Distortion [MW94b]. Rational [CD93]. Real [AT89, MN94, Bar88, Bar93, BH93]. recognition [Bon95]. record [IEE88]. recurrent [Har94]. Reducing [Sig96]. Reduction [SH94, WdJ94, Kom95b].

Region [TD95]. Region-Based [TD95]. Related [KCC87, Lep93]. Relating [vdW95]. Rendering [MD95]. Representation [KMK95, CL93]. resolution [KNH94]. results [CK96a]. Review [Jac93]. Rigidity [DBC93a, DBC93b]. Run [GK88a, GK88b]. Run-Length [GK88a, GK88b].

San [LUW95]. scale [Bog94a]. Scenes [KCC87, Pen84, Pen94]. Scheme [Gra92, RC94, LNF93]. Schemes [HM94, HS94, Hür95b, RFT94]. Science [PS88, PJS92a]. Scotland [Ano94]. Search [KNH94, Sau95a, TD95, HS93, RBM96, Sau95b]. searching [MW94a]. Segmentation [JR90]. Self [Dav95a, Dav95b, FRS94, Woh96, DV93, Ham95]. self-affine [DV93]. Self-Affinity [Woh96]. self-organizing [Ham95]. Self-Quantization [Dav95a]. Self-Quantized [Dav95b]. Self-VQ [FRS94]. September [Ano94]. Sequence [Reu93, LNF93]. Sequences [FRS94, MH91, MH92, VR94]. Set [CFMV92]. Sets [AT89, CFMV92]. several [Liu94]. side [SLTC94]. side-coupling [SLTC94]. Signal [IEE93, LUW95, Lun95, Vin93, Woh96, Ano94, ZCT94]. Similarity [HCF95]. Simple [PJS92b]. simplified [CK96b]. Single [VR94]. size [Sig96]. Snowbird [SC93, SC95]. Solution [HM90, Liu95b, HCF95]. solutions [Liu93b, Liu95a]. Solving [FV94a, FV94b]. Space [ØLR91, Woo94]. spanning [IEE88]. Speech [IEE93]. speed [BE95a]. Speeding [BE95b]. square [Sau96a]. state [CK96b]. Statistical [Hür95b]. Still [Bea90, HMS93, KP96, HS93]. Stochastic [Woh96]. storage [FL92]. Structural [CASH92, DV93]. Structure [TTMB96, SLTC94]. Structure-based [TTMB96]. Studies [DBC93a, DBC93b]. Subband [RC94, Sim95b]. Subimages [RC95]. Subtrees [Dav95a, Dav95b]. Symmetries [MW94b]. Synthesis [BJM^{+}88, SLN95]. System [JR90, MSK^{+}95, SLTC94, WY96]. system-based [WY96]. Systems [CFMV92, FV94a, FV94b, LY94a, LY94b, MH92, PD95, Reu93, Reu94a, Reu94b, Vin93, Har94].

target [EW96]. Technique [BSVN94, Jac90b]. Techniques [BDK92, GAH94, Lep93, Woo94]. technology [IEE88]. Tesselations [DBC93a, DBC93b]. textured [OCK95]. textured-image [OCK95]. their [Liu93b]. Theorem [ØBL^{+}94, Elt87]. Theoretical [FV94a]. Theory [Dav95a, Dav95b, Fis95, FV95, FGHS94, Jac92, KNH94, Reu93, Reu94a, Jac90a, Liu93a, BM92]. Theory-Based [KNH94]. Third [SC93]. Three [BV95]. Three-Dimensional [BV95]. Time [MN94, Sau94, WdJ94]. Transform [BSVN94, CZ97, KMK95, Mon93c, MW94b, Sim95a, vdW95, vdW96, Bon96, FSR94, HMS94, TW94, ZY94]. Transformation [BJ95, KNH94, BM92]. Transformations [Bog95b, Jac92, ØLR91, PJS92b, Reu93, Jac90a]. Transforms [FJB92, HH94, Mon93a, Mon93b, FV95, Hür93, WM94]. Tree [KNH94]. Triangular [DSC95]. Triangulation [DC94b, DC94a, DACB96]. Trondheim [Ano95]. Two [Bog94a, Hol91]. Two-Dimensional [Hol91]. Two-scale [Bog94a].

Unification [Liu94]. Unifying [BSVN94]. universe [IEE88]. USA [IEE93, IEE95, SC93, SC95]. Using [Bon95, CK93, FJB92, JR90, LY94a, LY94b, MH92, RC94, TD95, Bog95b, CK96b, FV94a, FV94b, GAH96, GIG95, Liu95a, MSK^{+}95, PD95, TW94, VR94, Woo94, ZP95, ZY94]. Utah [SC93, SC95].

Variable [MH91]. Vector [DACB96, HMS96a, KP94, KP96, BM95, ZP95]. via [AT89, KMK95, Sau95b]. Video [BV95, Bog94b, FRS94, LB94, MN94, WNM94, HB93]. vision [HCF95]. Visual [JR90, Che95, MSK^{+}95]. Volume [CHF94, CHF95]. VQ [FRS94, HMS96b]. VQ-enhanced [HMS96b].

Washington [IEE95]. Wavelet

[Dav95a, Dav95b, KMK95, LUW95, vdW96, CK96a, TW94]. Wavelet-Based [Dav95a, Dav95b]. wavelets [FSR94]. Way [BS88]. Weighted [CK93]. without [MW94a]. world [Bar88, Bar93, BH93]. Wyndham [IEE88].

Zooming [PD95].

References

Anonymous:1994:PEE

[Ano94] Anonymous, editor. Proceedings EUSIPCO’94: 7th European signal processing conference, September 1994, Edinburgh, Scotland. Loughborough University, Loughborough, UK, 1994. ISBN ???? LCCN ???? Three volumes.

Anonymous:1995:NAF

[Ano95] Anonymous, editor. NATO ASI on Fractal Image Encoding and Analysis: July 1995, Trondheim, Norway. ????, ????, 1995. ISBN ???? LCCN ????

Anson:1993:FIC

[Ans93] Louisa F. Anson. Fractal image compression. BYTE Magazine, pages 195–202, October 1993.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Abenda:1989:IPF

[AT89] S. Abenda and G. Turchetti. Inverse problems of fractal sets on the real line via the moment method. Nuovo Cimento, 104B: 213–227, 1989.

Abenda:1992:LMI

[ATD92] S. Abenda, G. Turchetti, and S. Demko. Local moments and inverse problem for fractal measures. Inverse Problems, 8:739– 750, 1992. CODEN INPEEY. ISSN 0266-5611 (print), 1361-6420 (electronic).

Barnsley:1988:FMR

[Bar88] Michael F. Barnsley. Fractal modelling of real world images. In Peitgen and Saupe [PS88], chapter 5, pages 219–242. ISBN 0-387-96608-0 (New York), 3-540-96608-0 (Berlin). LCCN QA614.86 S35 1988.

Barnsley:1993:FE

[Bar93] Michael F. Barnsley. Fractal modelling of real world images. In Fractals Everywhere, pages 219– 239. Academic Press, New York, NY, USA, 2nd edition, 1993. ISBN 0-12-079061-0.

Barthel:1995:ECF

[Bar95] Kai Uwe Barthel. Entropy constrained fractal image coding. In Anonymous [Ano95], page

????  ISBN ????  LCCN

????  URL ftp://ftp.
informatik.uni-freiburg.de/
papers/fractal/Bart95.ps.gz.

Belloulata:1996:FCM

[BBBCP96] Kamel Belloulata, Atilla M. Baskurt, Hugues Benoit-Cattin, and Remy Prost. Fractal coding of medical images. Proceedings of the SPIE — The International Society for Optical Engineering, 2707:598–609, 1996. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Barnsley:1986:CDF

[BD86] Michael F. Barnsley and Stephen G. Demko, editors. Chaotic Dynamics and Fractals, volume 2 of Notes and reports in mathematics in science and engineering. Academic Press, New York, NY, USA, 1 edition, 1986. ISBN 0-12-079060-2. xi + 292 pp. LCCN QA843 .C471 1986.

Bedford:1992:FIC

[BDK92] T. Bedford, F. M. Dekking, and M. S. Keane. Fractal image coding techniques and contraction operators. Nieuw Archief voor Wiskunde (Groningen), 10(3): 185–217, November 1992. CO-DEN NAWIA7. ISSN 0028-9825.

Bani-Eqbal:1995:ESF

[BE95a] Behnam Bani-Eqbal. Enhancing the speed of fractal image compression. Optical Engineering, 34(6):1705–1710, June 1995. CODEN OPEGAR. ISSN 0091-3286 (print), 1560-2303 (elec-

tronic). URL ftp://links.

uwaterloo.ca/pub/Fractals/
Papers/bani95.ps.gz.

Bani-Eqbal:1995:SFI

[BE95b] Behnam Bani-Eqbal. Speeding up fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2418:67–74, February 1995. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Beaumont:1990:ABB

[Bea90] J. M. Beaumont. Advances in block based fractal coding of still pictures. In Proceedings IEE Colloquium: The Application of Fractal Techniques in Image Processing, pages 3.1–3.6. ????, ????, December 1990.

Barnsley:1993:FIC

[BH93] Michael F. Barnsley and Lyman P. Hurd. Fractal modelling of real world images. In Fractal Image Compression, pages 219– 239. A. K. Peters, Ltd., Wellesley, MA, USA, 1993. ISBN 1-56881-000-8.

Boss:1995:ACI

[BJ95] R. D. Boss and E. W. Jacobs. Archetype classification in an iterated transformation image compression algorithm. In Yuval Fisher, editor, Fractal Image Compression: Theory and Application, chapter 4, pages 79– 90. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94211-4.

Barnsley:1988:HCI

[BJM^{+}88] Michael F. Barnsley, Arnaud Jacquin, François Malassenet, Laurie Reuter, and Alan D. Sloan. Harnessing chaos for image synthesis. Computer Graphics, 22(4):131–140, August 1988. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Bogdan:1992:KNN

[BM92] Alexandru Bogdan and H. E. Meadows. Kohonen neural network for image coding based on Iteration Transformation Theory. Proceedings of the SPIE — The International Society for Optical Engineering, 1766: 425–436, July 1992. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (elec-

tronic). URL ftp://ftp.ctr.
columbia.edu/CTR-Research/
advent/public/papers/92/bog92a.
ps.gz.

Bethel:1995:PIC

[BM95] David M. Bethel and D. M. Monro. Polynomial image coding with vector quantised compensation. In Proceedings ICASSP-95 (IEEE International Conference on Acoustics, Speech and Signal Processing), volume 4, pages 2499–2502. ????, ????, 1995. URL http:

//dmsun4.bath.ac.uk/papers/
icassp95bft.ps.gz.

Barahav:1993:HIF

[BMK93] Zachi Barahav, David Malah, and Ehud Karnin. Hierarchical interpretation of fractal image coding and its application to fast decoding. In IEEE [IEE93], pages 190–195. ISBN 0-7803-0947-2, 0-7803-0948-0 (paperback). LCCN TK 7882 S65 I16 1993. URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/BaMaKa93.ps.

Z. Five volumes. IEEE catalog number 93CH3252-4.

Barahav:1995:HIF

[BMK95] Zachi Barahav, David Malah, and Ehud Karnin. Hierarchical interpretation of fractal image coding and its applications. In Yuval Fisher, editor, Fractal Image Compression: Theory and Application, chapter 5, pages 91–117. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995.

Bogdan:1994:MFI

[Bog94a] Alexandru Bogdan. Multiscale fractal image coding and the two-scale difference equation. Technical Report CU/CTR/TR 358-94-05, Columbia University, New York, New York, March

1994. URL ftp://ftp.ctr.
columbia.edu/CTR-Research/
advent/public/papers/94/bog94a.
ps.Z.

Bogdan:1994:MII

[Bog94b] Alexandru Bogdan. Multiscale (inter/intra-frame) fractal video coding. In Proceedings ICIP-94 (IEEE International Conference on Image Processing), volume I, pages 760–764. ????, ????, November 1994. URL

                                        ftp://ftp.ctr.columbia.edu/
                                        CTR-Research/advent/public/
                                        papers/94/bog94b.ps.gz.

Bogdan:1995:FPA

[Bog95a] Alexandru Bogdan. The fractal pyramid with application to image coding. In Proceedings ICASSP-95 (IEEE International Conference on Acoustics, Speech and Signal Processing), volume 4, pages 2595–2598. ????, ????, May 1995. URL

                                        ftp://ftp.ctr.columbia.edu/
                                        CTR-Research/advent/public/
                                        papers/95/bog95a.ps.gz.

Bogdan:1995:ICU

[Bog95b] Alexandru Bogdan. Image Coding using Iterative Transformations with Applications to Image Communications. PhD thesis, Columbia University, New York, New York, 1995.

Bonneau:1995:UCO

[Bon95] Robert Bonneau. Using the centroid operator for faster multiresolution image compression and pattern recognition. Proceedings of the SPIE — The International Society for Optical Engineering, 2569:813–824, 1995. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Bonneau:1996:MTA

[Bon96] Robert Bonneau. Multiresolution transform and its application to image coding. Proceedings of the SPIE — The International Society for Optical Engineering, 2727:560–567, 1996. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Barnsley:1988:BWC

[BS88] Michael F. Barnsley and Alan D.

Sloan. A better way to compress images. BYTE Magazine, pages 215–223, January 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barthel:1994:NIC

[BSVN94] Kai Uwe Barthel, Jörg Schüttemeyer, Thomas Voyé, and Peter Noll. A new image coding technique unifying fractal and transform coding. In Proceedings ICIP-94 (IEEE International Conference on Image Processing), volume III, pages 112–116. ????, ????, November 1994. URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/BSVN94.ps.Z.

Barthel:1993:AFI

[BV93] Kai Uwe Barthel and Thomas Voyé. Adaptive fractal image coding in the frequency domain. In Proceedings of the International Workshop on Image Processing, Theory, Methodology, Systems and Applications. ????, ????, June 1993. URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/BaVo94.ps.Z.

Barthel:1994:AFI

[BV94] Kai Uwe Barthel and Thomas Voyé. Adaptive fractal image coding in the frequency domain.

In Proceedings of the International Workshop on Image Processing, volume XLV of Theory, Methodology, Systems and Applications, pages 33–38. ????, ????, June 1994. URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/BaVo94.ps.Z.

Barthel:1995:TDF

[BV95] Kai Uwe Barthel and Thomas Voyé. Three-dimensional fractal video coding. In Proceedings ICIP-95 (IEEE International Conference on Image Processing). ????, ????, 1995.

URL ftp://ftp.informatik.
uni-freiburg.de/papers/fractal/
BaVo95.ps.gz.

Barthel:1993:IFI

[BVN93] Kai Uwe Barthel, Thomas Voyé, and Peter Noll. Improved fractal image coding. In Proceedings PCS’93 (International Picture Coding Symposium), section 1.5, page 1.5. ????, ????, March 1993. URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/BaVoNo93.ps.
Z.

Cheong:1992:SED

[CASH92] Cha Keon Cheong, Kiyoharu Aizawa, Takahiro Saito, and Mitsutoshi Harori. Structural edge detection based on fractal analysis for image compression. In Proceedings of the IEEE International Symposium on Circuits and Systems, volume 5, pages 2461–2464. ????, ????, 1992.

Culik:1993:RAE

[CD93] Karel Culik and Simant Dube.

Rational and affine expression for image description. Discrete Applied Mathematics, 41(2):85– 120, 1993. CODEN DAMADU. ISSN 0166-218X (print), 1872-6771 (electronic).

Cabrelli:1992:IFS

[CFMV92] Carlos A. Cabrelli, Bruno Forte, Ursula M. Molter, and Edward R. Vrscay. Iterated fuzzy set systems: A new approach to the inverse problem for fractals and other sets. Journal of Mathematical Analysis and Applications, 171(1):79–100, November 1992. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Chen:1995:FIC

[Che95] Jer-Sen Chen. Fractal image compression based on visual perception. Proceedings of the SPIE — The International Society for Optical Engineering, 2411: 92–99, 1995. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Cochran:1994:FVC

[CHF94] Wayne O. Cochran, John C.

Hart, and Patrick J. Flynn. Fractal volume compression. Internal report, Washington State University, School of EECS, Pullman, WA, 1994.

Cochran:1995:FVC

[CHF95] Wayne O. Cochran, John C.

Hart, and Patrick J. Flynn. Fractal volume compression. IEEE transactions on visualization and computer graphics, October 1995. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Culik:1993:ICU

[CK93] Karel Culik and J. Kari. Image compression using weighted finite automata. Computers and Graphics, 17(3):305–313, 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Caso:1996:NRF

[CK96a] Gregory Caso and C.-C. J. Kuo. New results for fractal/wavelet image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2727:536–547, 1996. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Chang:1996:FBC

[CK96b] Hsuan T. Chang and Chung Jung Kuo. Fractal block coding using a simplified finite-state algorithm. Proceedings of the SPIE — The International Society for Optical Engineering, 2501:536–544, 1996. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Clarke:1993:FIR

[CL93] R. J. Clarke and L. M. Linnett. Fractals and image representation. Electronics and Communication Engineering Journal, 5(4):

233–239, August 1993. CODEN ECEJE9. ISSN 0954-0695.

Caso:1996:FMF

[COK96] Gregory Caso, Pere Obrador, and C.-C. J. Kuo. Fast methods for fractal image encoding. Proceedings of the SPIE — The International Society for Optical Engineering, 2501:583–594, 1996. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Cheng:1997:MAF

[CZ97] Bing Cheng and Xiaokun Zhu.

Multiresolution approximation of fractal transform. Submitted to Signal Processing., 1997. CO-DEN SPRODR. ISSN 0165-1684 (print), 1872-7557 (electronic). Preprint.

Davoine:1996:FIC

[DACB96] Franck Davoine, Marc Antonini, Jean-Marc Chassery, and Michel Barlaud. Fractal image compression based on Delaunay triangulation and vector quantization. IEEE Transactions on Image Processing, 5(2):338–346, February 1996. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Davis:1995:ASQ

[Dav95a] Geoffrey Davis. Adaptive selfquantization of wavelet subtrees: A wavelet-based theory of fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2569:294–307, July 1995. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (elec-

tronic). URL http://www.cs.
dartmouth.edu/~gdavis/papers/
spie_sandiego.ps.g.

Davis:1995:SQW

[Dav95b] Geoffrey Davis. Self-quantized wavelet subtrees: A waveletbased theory of fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2491:141–152, April 1995. CODEN PSISDG. ISBN 0-8194-1844-7. ISSN 0277-786X (print), 1996-756X (elec-

tronic). URL http://www.cs.
dartmouth.edu/~gdavis/spie_
fractal.ps.gz.

Davis:1996:IIM

[Dav96] Geoffrey Davis. Implicit image models in image fractal compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2569:88–97, 1996. CODEN PSISDG. ISBN 0-8194-2213-4. ISSN 0277-786X (print), 1996-756X (electronic).

Davoine:1993:IWI

[DBC93a] F. Davoine, E. Bertin, and J-M. Chassery. From rigidity to adaptive tesselations for fractal image compression: Comparative studies. In IEEE 8th Workshop on Image and Multidimensional Signal Processing. ????, ????, September 1993.

URL ftp://ftp.informatik.
uni-freiburg.de/documents/
papers/fractal/DaBeCh93.ps.
Z.

Davoine:1993:RAT

[DBC93b] Franck Davoine, Etienne Bertin, and Jean-Marc Chassery. From rigidity to adaptive tesselations for fractal image compression: Comparative studies. In Proceedings of the 8th IEEE Workshop on Image and Multidimensional Signal Processing, pages 56– 57. ????, ????, September 1993.

URL ftp://ftp.informatik.
uni-freiburg.de/documents/
papers/fractal/DaBeCh93.ps.
Z.

Davoine:1994:ICP

[DC94a] F. Davoine and J-M. Chassery.

Adaptive delaunay triangulation for attractor image coding. In 12th International Conference on Pattern Recognition. ????, ????, October 1994. URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/DaCh94.ps.Z.

to appear.

Davoine:1994:ADT

[DC94b] Franck Davoine and Jean-Marc Chassery. Adaptive Delaunay triangulation for attractor image coding. In Proceedings of the 12th International Conference on Pattern Recognition, pages 801–803. ????, ????, October 1994. URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/DaCh94.ps.Z.

Dettmer:1992:FFF

[Det92] R. Dettmer. Form and function:

Fractal-based image compression. IEE Review, 38(9):323–327, 1992. CODEN IEREEF. ISSN 0953-5683 (or 0013-5127??).

Devaney:1989:CFM

[DK89] Robert L. Devaney and Linda Keen, editors. Chaos and Fractals (The Mathematics Behind the Computer Graphics), volume 39 of Proceedings of Symposia in Applied Mathematics. Amer. Math. Soc., Providence, RI, USA, 1989. ISBN 0-8218-0137-6. ix + 148 pp. LCCN QA1 .S93 v.39.

Davoine:1995:MTQ

[DSC95] Franck Davoine, John Svensson, and Jean-Marc Chassery. A mixed triangular and quadrilateral partition for fractal image coding. In Proceedings ICIP-95 (IEEE International Conference on Image Processing), volume III, pages 284–287. ????, ????, October 1995.

Dudbridge:1994:FIC

[Dud94] Frank Dudbridge. Fast image coding by a hierarchical fractal construction. Unknown, 1994. Submitted.

Domaszewicz:1993:SLS

[DV93] Jaroslaw Domaszewicz and Vinay A. Vaishampayan. Structural limitations of self-affine and partially self-affine fractal compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2094:1498– 1507, 1993. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Domaszewicz:1994:ICC

[DV94] Jaroslaw Domaszewicz and Vinay A.

Vaishampayan. Iterative collage coding for fractal compression. In Proceedings ICIP-94 (IEEE International Conference on Image Processing), volume III, pages 127–131. ????, ????, November 1994.

Elton:1987:ETI

[Elt87] John H. Elton. An ergodic theorem for iterated maps. Ergodic Theory and Dynamical Systems, 7:481–488, 1987. ISSN 0143-3857 (print), 1469-4417 (electronic).

Ewing:1996:CJF

[EW96] Gary J. Ewing and Christopher J. Woodruff. Comparison of JPEG and fractal-based image compression on target acquisition by human observers. Optical Engineering, 35(1):284–288, January 1996. CODEN OPE-GAR. ISSN 0091-3286 (print), 1560-2303 (electronic).

Frigaard:1994:ICB

[FGHS94] Carsten Frigaard, Jess Gade, Thomas Hemmingsen, and Torben Sand. Image compression based on a fractal theory. Internal Report S701, Institute for Electronic Systems, Aalborg University, Aalborg, Denmark, 1994. URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/FGHS94.ps.gz.

Fisher:1992:DFI

[Fis92a] Yuval Fisher. A discussion of fractal image compression. In Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe, editors, Chaos and Fractals: New Frontiers of Science, app. Appendix A, pages 903–919. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97903-4.

Fisher:1992:FICa

[Fis92b] Yuval Fisher. Fractal image compression. Technical Report 12, Department of Mathematics, Technion Israel Institute of Technology, 1992. 7.1–7.19 pp.

URL ftp://lyapunov.ucsd.
edu/pub/inls-ucsd/fractal_

paper.ps.Z. SIGGRAPH ‘92 COURSE NOTES.

Fisher:1995:FIC

[Fis95] Yuval Fisher, editor. Fractal Image Compression: Theory and Application. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94211-4.

Fisher:1992:FICc

[FJB92] Y. Fisher, E. W. Jacobs, and R. D. Boss. Fractal image compression using iterated transforms. In James A. Storer, editor, Image and Text Compression, chapter 2, pages 35– 61. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-9243-4.

Fisher:1992:FICb

[FL92] Yuval Fisher and Albert F. Lawrence. Fractal image compression for mass storage applications. Proceedings of the SPIE — The International Society for Optical Engineering, 1662:244– 255, 1992. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Fisher:1994:FSV

[FRS94] Yuval Fisher, Daniel N. Rogovin, and Tsae-Pyng J. Shen. Fractal (self-VQ) encoding of video sequences. Proceedings of the SPIE — The International Society for Optical Engineering, 2308:1359–1370, September 28– 29, 1994. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic). URL http:

//inls.ucsd.edu/y/Fractals/
Video/spie.ps.Z.

Fisher:1994:CFM

[FSR94] Y. Fisher, T. P. Shen, and D. Rogovin. A comparison of fractal methods with discrete cosine transform (DCT) and wavelets. Proceedings of the SPIE — The International Society for Optical Engineering, 2304-16:132– 143, July 28–29, 1994. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

URL http://inls.ucsd.edu/y/
Fractals/spie.ps.

Forte:1994:SIPa

[FV94a] Bruno Forte and Edward Vrscay.

Solving the inverse problem for function/image approximation using iterated function systems: I theoretical basis. Fractals, 2(3):325–334, 1994. CODEN FRACEG. ISSN 0218-348X.

Forte:1994:SIPb

[FV94b] Bruno Forte and Edward Vrscay. Solving the inverse problem for function/image approximation using iterated function systems: II algorithms and computations. Fractals, 2(3):335–346, 1994. CO-DEN FRACEG. ISSN 0218-348X.

Forte:1995:TGF

[FV95] Bruno Forte and Edward Vrscay. Theory of generalized fractal transforms. In Anonymous [Ano95], page ???? ISBN ???? LCCN ???? Gharavi-Alkhansari:1993:FBI

[GAH93] Mohammad Gharavi-Alkhansari and Thomas S. Huang. A fractalbased image block-coding algorithm. In Proceedings PCS’93 (International Picture Coding Symposium), page 1.7. ????, ????, March 1993.

Gharavi-Alkhansari:1994:FBT

[GAH94] Mohammad Gharavi-Alkhansari and Thomas S. Huang. Fractalbased techniques for a generalized image coding method. In Proceedings ICIP-94 (IEEE International Conference on Image Processing), volume III, pages 122–126. ????, ????, November 1994. Gharavi-Alkhansari:1996:FIC

[GAH96] Mohammad Gharavi-Alkhansari and Thomas S. Huang. Fractal image coding using ratedistortion optimized matching pursuit. Proceedings of the SPIE — The International Society for Optical Engineering, 2727:1386– 1393, 1996. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Gotting:1995:FIC

[GIG95] Detlef Götting, Achim Ibenthal, and Rolf-Rainer Grigat. Fractal image coding and magnification using invariant features. In Anonymous [Ano95], page

????  ISBN ????  LCCN

????  URL ftp://ftp.
informatik.uni-freiburg.de/
papers/fractal/GoIbGr95.ps.
gz.

Goel:1988:DCA

[GK88a] B. Goel and S. Kwatra. A data compression algorithm for color images based on run-length coding and fractal geometry. In IEEE [IEE88], pages 1253–1256. ISBN ???? LCCN TK5101.A1 I34 1988. IEEE catalog number 88CH2538-7.

Goel:1988:IIC

[GK88b] B. Goel and S. Kwatra. A data compression algorithm for color images based on run-length coding and fractal geometry. In IEEE [IEE88], pages 1253–1256. ISBN ???? LCCN TK5101.A1 I34 1988. IEEE catalog number 88CH2538-7.

Graf:1992:BSF

[Gra92] S. Graf. Barnsley’s scheme for the fractal encoding of images. Journal of Complexity, 8:72–78, 1992.

CODEN JOCOEH. ISSN 0885-064X (print), 1090-2708 (electronic).

Hamzaoui:1995:CCS

[Ham95] Raouf Hamzaoui. Codebook clustering by self-organizing maps for fractal image compression. In Anonymous [Ano95], page

????  ISBN ????  LCCN

                                        ????  URL ftp://ftp.
                                        informatik.uni-freiburg.de/
                                        papers/fractal/Hamz95.ps.gz.

Hart:1994:FIC

[Har94] John C. Hart. Fractal image compression and the inverse problem of recurrent iterated function systems. IEEE Computer Graphics and Applications, July 1994. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hurtgen:1993:FAL

[HB93] Bernd Hürtgen and Peter Büttgen. Fractal approach to low rate video coding. Proceedings of the SPIE — The International Society for Optical Engineering, 2094: 120–131, November 1993. CO-DEN PSISDG. ISBN 0-8194-1369-0. ISSN 0277-786X (print), 1996-756X (electronic).

Hart:1995:SHC

[HCF95] John C. Hart, Wayne O. Cochran, and Patrick J. Flynn. Similarity hashing: A computer vision solution to the inverse problem of linear fractals. In Anonymous [Ano95], page ???? ISBN ???? LCCN ????

Hurtgen:1994:CFT

[HH94] Bernd Hürtgen and Thomas Hain. On the convergence of fractal transforms. In Proceedings ICASSP-94 (IEEE International Conference on Acoustics, Speech and Signal Processing), volume 5, pages 561–564. ????, ????, April 1994. ISBN 0-7803-1775-0 (softbound), 0-7803-1776-9 (casebound), 0-7803-1777-7 (microfiche). LCCN TK 7882 S65 I16 1994.

Handy:1990:IPF

[HM90] C. Handy and Giorgio Mantica.

Inverse problems in fractal construction: Moment method solution. Physica. D, Nonlinear phenomena, 43:17–36, 1990. CO-DEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic).

Hurtgen:1994:MFC

[HM94] Bernd Hürtgen and Frank Müller. Modelling of fractal coding schemes. In Anonymous [Ano94], pages 600–603. ISBN ???? LCCN ???? Three volumes.

Hurtgen:1993:AFC

[HMS93] Bernd Hürtgen, Frank Müller, and Christoph Stiller. Adaptive fractal coding of still pictures. In Proceedings PCS’93 (International Picture Coding Symposium), page 1.8. ????, ????, March 1993.

Hurtgen:1994:FTC

[HMS94] Bernd Hürtgen, Paul Mols, and Stephan F. Simon. Fractal transform coding of color images. Proceedings of the SPIE — The International Society for Optical Engineering, 2308:1683–1691, July 1994. CODEN PSISDG. ISBN 0-8194-1638-X. ISSN 0277-786X (print), 1996-756X (electronic).

Hamzaoui:1996:EFI

[HMS96a] Raouf Hamzaoui, Martin Müller, and Dietmar Saupe. Enhancing fractal image compression with vector quantization. In IEEE Digital Signal Processing Workshop. ????, ????, September 1996. URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/HaMuSa96b.ps.

gz. Preprint.

Hamzaoui:1996:VEF

[HMS96b] Raouf Hamzaoui, Martin Müller, and Dietmar Saupe. VQenhanced fractal image compression. In Proceedings ICIP-96 (IEEE International Conference on Image Processing), volume I, pages 153–156. ????, ????, September 1996. URL ftp://

ftp.informatik.uni-freiburg.
de/papers/fractal/HaMuSa96a.
ps.gz.

Hollatz:1991:DIC

[Hol91]  Scott A. Hollatz.  Digital

image compression with twodimensional affine fractal interpolation functions. Technical Report 91-2, Department of Mathematics and Statistics, University of Minnesota - Duluth, 1991.

URL ftp://hp.uwsuper.edu/
pub/fic.tar.Z.

Hurtgen:1993:FHC

[HS93] Bernd Hürtgen and Christoph Stiller. Fast hierarchical codebook search for fractal coding of still images. Proceedings of the SPIE — The International Society for Optical Engineering, 1977: 397–408, April 1993. CODEN PSISDG. ISBN 0-8194-1223-6. ISSN 0277-786X (print), 1996-756X (electronic).

Hurtgen:1994:PCF

[HS94] Bernd Hürtgen and Stephan F.

Simon. On the problem of convergence in fractal coding schemes. In Proceedings ICIP-94 (IEEE International Conference on Image Processing), volume III, pages 103–106. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, November 1994. ISBN 0-8186-6952-7 (case), 0-8186-6950-0 (paper), 0-8186-6951-9 (microfiche). LCCN TA 1637 I25 1994.

Hurtgen:1993:CFT

[Hür93] Bernd Hürtgen. Contractivity of fractal transforms for image coding. Electronics Letters, 29(20): 1749–1750, September 1993. CO-DEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Hurtgen:1995:PBF

[Hür95a] Bernd Hürtgen. Performance bounds for fractal coding. In Proceedings ICASSP-95 (IEEE International Conference on Acoustics, Speech and Signal Processing), volume 4, pages 2563–2566. ????, ????, May 1995. ISBN 0-7803-2431-5 (softbound), 0-7803-2432-3 (casebound), 0-7803-2433-1 (microfiche), 0-7803-2562-1 (CD-ROM). LCCN TK7882.S65 .I16 1995.

Hurtgen:1995:SEF

[Hür95b] Bernd Hürtgen. Statistical evaluation of fractal coding schemes. In Proceedings ICIP-95 (IEEE International Conference on Image Processing), volume III, pages 280–283. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 1995. LCCN TK8315.I222 1995.

IEEE:1988:DTS

[IEE88]  IEEE, editor.  Digital tech-

nology — spanning the universe: IEEE International Conference on Communications ’88: Philadelphia, PA, June 12–15, 1988, Wyndham Franklin Plaza Hotel: conference record. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN ???? LCCN TK5101.A1 I34 1988. IEEE catalog number 88CH2538-7.

IEEE:1993:IIC

[IEE93] IEEE, editor. 1993 IEEE International Conference on Acoustics, Speech, and Signal Processing, April 27–30, 1993, Minneapolis Convention Center, Minneapolis, Minnesota, USA. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-0947-2, 0-7803-0948-0 (paperback). LCCN TK 7882 S65 I16 1993. Five volumes. IEEE catalog number 93CH3252-4.

IEEE:1995:IPI

[IEE95] IEEE, editor. ICIP’95: Proceedings, International Conference on Image Processing: October 23–26, 1995, Washington, DC, USA. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-3122-2 (casebound), 0-8186-7310-9 (softbound), 0-7803-3123-0 (microfiche), 0-7803-2749-7 (CD-ROM). LCCN TK8315.I222 1995. Three volumes. IEEE catalog number 95CB35819.

Jacquin:1990:FIC

[Jac90a] Arnaud E. Jacquin. Fractal image coding based on a theory of iterated contractive image transformations. Proceedings of the SPIE — The International Society for Optical Engineering, 1360: 227–239, October 1990. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Jacquin:1990:NFB

[Jac90b] Arnaud E. Jacquin. A novel fractal block-coding technique for digital images. In Proceedings ICASSP-90 (IEEE International Conference on Acoustics, Speech and Signal Processing), volume 4, pages 2225–2228. ????, ????, April 1990.

Jacquin:1992:ICB

[Jac92] Arnaud E. Jacquin. Image coding based on a fractal theory of iterated contractive image transformations. IEEE Transactions on Image Processing, 1(1):18–30, January 1992. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Jacquin:1993:FIC

[Jac93] Arnaud E. Jacquin. Fractal image coding: A review. Proceedings of the IEEE, 81(10): 1451–1465, October 1993. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Jang:1990:SBI

[JR90] Jongwhan Jang and Sarah A. Rajala. Segmentation based image coding using fractals and the human visual system. In Proceedings ICASSP-90 (IEEE International Conference on Acoustics, Speech and Signal Processing), volume 4, pages 1957–1960. ????, ????, April 1990.

Keller:1987:CNS

[KCC87] James M. Keller, Richard M. Crownover, and Robert Yu Chen. Characteristics of natural scenes related to the fractal dimension. IEEE Transactions on Pattern Analysis and Machine Intelligence, PAMI-9(5):621–627, September 1987. CODEN IT-PIDJ. ISSN 0162-8828.

Kim:1995:NFI

[KKL95] Chang-Su Kim, Rin-Chul Kim, and Sang-Uk Lee. Novel fractal image compression method with non-iterative decoder. In IEEE [IEE95], pages 268–271. ISBN 0-7803-3122-2 (casebound), 0-8186-7310-9 (softbound), 0-7803-3123-0 (microfiche), 0-7803-2749-7 (CD-ROM). LCCN TK8315.I222 1995. Three volumes. IEEE catalog number 95CB35819.

Krupnik:1995:FRI

[KMK95] Hagai Krupnik, David Malah, and Ehud Karnin. Fractal representation of images via the discrete wavelet transform. In Proceedings of the 18th IEEE Convention of Electrical and Electronics Engineers in Israel, page 2.2.2. ????, ????, March 1995.

URL http://inls.ucsd.edu/y/
Fractals/hagai_paper.ps.gz.

Kawamata:1994:MRT

[KNH94] Masayuki Kawamata, Masaki Nagahisa, and Tatsuo Higuchi. Multi-resolution tree search for iterated transformation theorybased coding. In Proceedings ICIP-94 (IEEE International Conference on Image Processing), volume III, pages 137–141. ????, ????, November 1994.

Kominek:1995:AFF

[Kom95a] John Kominek. Algorithm for fast fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2419:296– 305, February 1995. CODEN PSISDG. ISBN 0-8194-1766-1. ISSN 0277-786X (print), 1996-756X (electronic).

Kominek:1995:CRF

[Kom95b] John Kominek. Codebook reduction in fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2669:33–41, 1995. CODEN PSISDG. ISBN 0-8194-2043-3. ISSN 0277-786X (print), 1996-756X (electronic).

Kominek:1995:CFE

[Kom95c] John Kominek. Convergence of fractal encoded images. In Storer and Cohn [SC95], pages 242–251. ISBN 0-8186-7012-6. ISSN 1068-0314. LCCN QA76.9.D33 D37 1995.

Kim:1994:ICB

[KP94] Kwon Kim and Rae-Hong Park. Image coding based on fractal approximation and vector quantization. In Proceedings ICIP-94 (IEEE International Conference on Image Processing), volume III, pages 132–136. ????, ????, November 1994.

Kim:1996:SIC

[KP96] Kwon Kim and Rae-Hong Park. Still image coding based on vector quantization and fractal approximation. IEEE Transactions on Image Processing, 5(4): 587–597, April 1996. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Kuroda:1996:FBM

[KPY96] Hideo Kuroda, Dan C. Popescu, and Hong Yan. Fast block matching method for image data compression based on fractal models. Proceedings of the SPIE — The International Society for Optical Engineering, 2501:1257– 1266, 1996. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Lazar:1994:FBC

[LB94] M. S. Lazar and L. T. Bruton. Fractal block coding of digital video. IEEE Transactions on Circuits and Systems for Video Technology, 4(3):297–308, June 1994. CODEN ITCTEM. ISSN 1051-8215 (print), 1558-2205 (electronic). URL http://

www-mddsp.enel.ucalgary.ca/
Papers/msl94-2.ps.gz.

Lepsoy:1993:AIC

[Lep93] Skjalg Lepsøy. Attractor Image Compression - Fast Algorithms and Comparisons to Related Techniques. PhD thesis, The Norwegian Institute of Technology, Trondheim, Norway, June 1993.

Lin:1994:FIC

[Lin94] David W. Lin. Fractal image coding as generalized predictive coding. In Proceedings ICIP-94 (IEEE International Conference on Image Processing), volume III, pages 117–121. ????, ????, November 1994.

Liu:1993:EFT

[Liu93a] Ying Liu. Extensions of fractal theory. Proceedings of the SPIE — The International Society for Optical Engineering, 1966:255– 268, 1993. CODEN PSISDG.

ISSN 0277-786X (print), 1996-756X (electronic).

Liu:1993:FET

[Liu93b] Ying Liu. Fractal equations and their solutions. Proceedings of the SPIE — The International Society for Optical Engineering, 1904: 52–68, 1993. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Liu:1994:USI

[Liu94] Ying Liu. Unification of several image compression methods. Proceedings of the SPIE — The International Society for Optical Engineering, 2238:92–106, 1994. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Liu:1995:IBC

[Liu95a] Ying Liu. Image block coding using exact solutions of fractal equations. Proceedings of the SPIE — The International Society for Optical Engineering, 2488:135–149, 1995. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Liu:1995:SFE

[Liu95b] Ying Liu. Solution of fractal equation. Proceedings of the SPIE — The International Society for Optical Engineering, 2418: 48–59, 1995. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Loew:1996:APM

[LLP96] Murray H. Loew, Dunling Li, and Raymond L. Pickholtz. Adaptive PIFS model in fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2707:284– 293, 1996. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Li:1993:FBI

[LNF93] Haibo Li, Mirek Novak, and Robert Forchheimer. Fractalbased image sequence compression scheme. Optical Engineering, 32(7):1588–1595, July 1993. CO-DEN OPEGAR. ISSN 0091-3286 (print), 1560-2303 (electronic).

Lepsoy:1995:FAI

[LØ95] Skjalg Lepsøy and Geir E. Øien.

Fast attractor image encoding by adaptive codebook clustering. In Yuval Fisher, editor, Fractal Image Compression: Theory and Application, chapter 9, pages 177–197. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94211-4.

Lundheim:1995:DFF

[Lun95] Lars M. Lundheim. A discrete framework for fractal signal modeling. In Yuval Fisher, editor, Fractal Image Compression: Theory and Application, chapter 7, pages 137–151. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94211-4.

Laine:1995:SQW

[LUW95] Andrew F. Laine, Michael A. Unser, and Mladen V. Wickerhauser, editors. Wavelet Applications in Signal and Image Processing III: 12–14 July, 1995, San Diego, California, volume 2569 of Proceedings of the SPIE — The International Society for Optical Engineering. Society of Photo-optical Instrumentation Engineers (SPIE), Bellingham, WA, USA, 1995. ISBN 0-8194-1928-1. LCCN TS510.S63 no.2569.

Lin:1995:PAF

[LV95] H. Lin and A. N. Venetsanopoulos. A pyramid algorithm for fast fractal image compression. In Proceedings ICIP-95 (IEEE International Conference on Image Processing), volume III, pages 596–599. ????, ????, October 1995.

Lin:1996:PLF

[LV96] Huawu Lin and Anastasios N. Venetsanopoulos. Perceptually lossless fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2727:1394– 1399, 1996. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Lu:1994:ICUa

[LY94a] G. Lu and T. L. Yew. Image compression using quadtree partitioned iterated function systems. Electronics Letters, 30 (1):23–24, January 1994. CO-DEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Lu:1994:ICUb

[LY94b] Guojun Lu and Toon-Lin Yew.

Image compression using partitioned iterated function systems. Proceedings of the SPIE — The International Society for Optical Engineering, 2186:122–133, 1994. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Monro:1992:FAI

[MD92a] Donald M. Monro and Frank Dudbridge. Fractal approximation of image blocks. In Proceedings ICASSP-92 (IEEE International Conference on Acoustics, Speech and Signal Processing), volume 3, pages 485–488. ????, ????, March 1992. CO-DEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

URL http://dmsun2.bath.ac.
uk:8080/people/papers/assp92.
ps.gz.

Monro:1992:FBC

[MD92b] Donald M. Monro and Frank Dudbridge. Fractal block coding of images. Electronics Letters, 28 (11):1053–1055, May 1992. CO-DEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Monro:1995:RAD

[MD95] Donald M. Monro and Frank Dudbridge. Rendering algorithms for deterministic fractals. IEEE Computer Graphics and Applications, 15(1):32–41, January 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Mazel:1991:HVF

[MH91] David S. Mazel and Monson H. Hayes. Hidden-variable fractal interpolation of discrete sequences. Electronics Letters, pages 3393– 3396, May 1991. CODEN EL-LEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Mazel:1992:UIF

[MH92] David S. Mazel and Monson H. Hayes. Using iterated function systems to model discrete sequences. IEEE Transactions on Signal Processing, 40(7):1724– 1734, July 1992. CODEN IT-PRED. ISSN 1053-587X (print), 1941-0476 (electronic).

Monro:1994:RTF

[MN94] Donald M. Monro and Jeremy A. Nicholls. Real time fractal video for personal communications. Fractals, 2(6):39–42, 1994. CODEN FRACEG. ISSN 0218-348X.

Monro:1993:CFT

[Mon93a] Donald M. Monro. Class of fractal transforms. Electronics Letters, 29(4):362–363, February 1993. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Monro:1993:GFT

[Mon93b] Donald M. Monro. Generalized fractal transforms: Complexity issues. In Storer and Cohn [SC93], pages 254–261. ISBN 0-8186-3392-1, 0-8186-3391-3. ISSN 1068-0314. LCCN QA76.9.D33 D38 1993. IEEE catalog number 93TH0536-3. IEEE Computer Society Press order number 3392-02.

Monro:1993:HFT

[Mon93c] Donald M. Monro. A hybrid fractal transform. Electronics Letters, 5:169–172, 1993. CO-DEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

URL http://dmsun2.bath.ac.
uk:8080/people/papers/assp93.
ps.gz.

Moon:1995:FIC

[MSK^{+}95] Yong H. Moon, Kyung S. Son, Hyung S. Kim, Yoon S. Kim, and Jae H. Kim. Fractal image compression using human visual system. Proceedings of the SPIE — The International Society for Optical Engineering, 2418: 60–66, 1995. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Monro:1994:FIC

[MW94a] Donald M. Monro and Stuart J.

Woolley. Fractal image compression without searching. In Proceedings ICASSP-94 (IEEE International Conference on Acoustics, Speech and Signal Processing), volume 5, pages 557– 560. ????, ????, April 1994.

URL http://dmsun2.bath.ac.
uk:8080/people/papers/bft94.
ps.gz.

Monro:1994:RDF

[MW94b] Donald M. Monro and Stuart J. Woolley. Rate/distortion in fractal compression: Order of transform and block symmetries. In Proceedings ISSIPNN’94 (International Symposium on Speech, Image Processing and Neural Networks), volume 1, pages 168– 171. ????, ????, April 1994. ISBN 0-7803-1865-X.

Novak:1993:ACI

[Nov93a] Mirek Novak. Attractor coding of images. In Proceedings PCS’93 (International Picture Coding Symposium), page 15.6. ????, ????, March 1993.

                                        URL ftp://ftp.informatik.
                                        uni-freiburg.de/papers/fractal/
                                        Nova93b.ps.Z.

Novak:1993:ACIb

[Nov93b] Mirek Novak. Attractor coding of images. In Proceedings from Picture Coding Symposium (PCS). ????, ????, March 1993. URL ftp://ftp.

                                        informatik.uni-freiburg.de/
                                        papers/fractal/Nova93b.ps.
                                        Z.

Novak:1993:ACIa

[Nov93c] Miroslav Novak. Attractor Coding of Images. PhD thesis, Department of Electrical Engineering, Linköping University, Linköping, Sweden, May 1993. URL ftp://ftp.

                                        informatik.uni-freiburg.de/
                                        papers/fractal/Nova93a.ps.
                                        Z.

Oien:1994:NIC

[ØBL^{+}94] Geir E. Øien, Zachi Baharav, Skjalg Lepsøy, Ehud Karnin, and David Malah. A new improved collage theorem with applications to multiresolution fractal image coding. In Proceedings ICASSP-94 (IEEE International Conference on Acoustics, Speech and Signal Processing), volume 5, pages 565–568. ????, ????, April 1994. URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/OBLMK94.ps.

Z. submitted.

Obrador:1995:FBM

[OCK95] Pere Obrador, Gregory Caso, and C.-C. J. Kuo. Fractal-based method for textured-image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2418: 36–47, 1995. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Oien:1994:PQF

[Øie94] Geir E. Øien. Parameter quantization in fractal image coding. In Proceedings ICIP-94 (IEEE International Conference on Image Processing), volume III, pages 142–146. ????, ????, November 1994.

Oien:1995:CFI

[ØL95] Geir E. Øien and Skjalg Lepsøy.

A class of fractal image coders with fast decoder convergence. In Yuval Fisher, editor, Fractal Image Compression: Theory and Application, chapter 8, pages 153–175. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94211-4.

Oien:1991:IPS

[ØLR91] Geir E. Øien, Skjalg Lepsøy, and Tor A. Ramstad. An inner product space approach to image coding by contractive transformations. In Proceedings ICASSP-91 (IEEE International Conference on Acoustics, Speech and Signal Processing), pages 2773– 2776. ????, ????, May 1991.

Polidori:1995:ZUI

[PD95] Eric Polidori and Jean-Luc Dugelay. Zooming using iterated function systems. In Anonymous [Ano95], page ???? ISBN ???? LCCN ????

Popescu:1997:NMF

[PDY97] Dan C. Popescu, Alex Dimca, and Hong Yan. A nonlinear model for fractal image coding. IEEE Transactions on Image Processing, 6(3):373–382, March 1997. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Pentland:1984:FBD

[Pen84]  Alex P. Pentland.  Fractal-

based description of natural scenes. IEEE Transactions on Pattern Analysis and Machine Intelligence, PAMI-6(6):661–674, November 1984. CODEN IT-PIDJ. ISSN 0162-8828.

Pentland:1994:FBD

[Pen94]  Alex P. Pentland.  Fractal-

based description of natural scenes. IEEE Transactions on Pattern Analysis and Machine Intelligence, PAMI-6(6):661–674, November 1994. CODEN IT-PIDJ. ISSN 0162-8828.

Pentland:1991:PAF

[PH91] Alexander P. Pentland and Bradley Horowitz. Practical approach to fractal-based image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 1605:467–474, 1991. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Peitgen:1992:CFN

[PJS92a] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Chaos and fractals: New frontiers of science. In Chaos and Fractals: New Frontiers of Science, chapter 5, pages 229–296. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97903-4. LCCN QA614.86 .P43 1992.

Peitgen:1992:EIS

[PJS92b] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Encoding images by simple transformations. In Chaos and Fractals: New Frontiers of Science, chapter 5, pages 229–296. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97903-4.

Peitgen:1988:SFI

[PS88] Heinz-Otto Peitgen and Dietmar Saupe, editors. The Science of Fractal Images. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96608-0 (New York), 3-540-96608-0 (Berlin). xiii + 312 pp. LCCN QA614.86 S35 1988.

Redmill:1996:GAF

[RBM96] David W. Redmill, David R. Bull, and Ralph R. Martin. Genetic algorithms for fast search in fractal image coding. Proceedings of the SPIE — The International Society for Optical Engineering, 2727:1367–1376, 1996. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Rinaldo:1994:ICS

[RC94] Roberto Rinaldo and Giancarlo Calvagno. An image coding scheme using block prediction of the pyramid subband decomposition. In Proceedings ICIP-94 (IEEE International Conference on Image Processing), volume II, pages 878–882. ????, ????, November 1994. URL ftp://ftp.

                                        informatik.uni-freiburg.de/
                                        papers/fractal/RiCa94.ps.gz.

Rinaldo:1995:ICB

[RC95] Roberto Rinaldo and Giancarlo Calvagno. Image coding by block prediction of multiresolution subimages. IEEE Transactions on Image Processing, 4 (7):909–920, July 1995. CO-DEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic). URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/RiCa95.ps.gz.

to appear.

Reusens:1993:SCB

[Reu93] Emmanuel Reusens. Sequence coding based on the fractal theory of iterated transformations systems. Proceedings of the SPIE — The International Society for Optical Engineering, 2094: 132–140, November 1993. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Reusens:1994:OAP

[Reu94a] Emmanuel Reusens. Overlapped adaptive partitioning for image coding based on the theory of iterated functions systems. In Proceedings ICASSP-94 (IEEE International Conference on Acoustics, Speech and Signal Processing), volume 5, pages 569–572. ????, ????, April 1994.

Reusens:1994:PCI

[Reu94b] Emmanuel Reusens. Partitioning complexity issue for iterated functions systems based image coding. In Anonymous [Ano94], pages 171–174. ISBN ???? LCCN ???? Three volumes.

Ratnakar:1994:FBH

[RFT94] Viresh Ratnakar, Ephraim Feig, and Prasoon Tiwari. Fractalbased hybrid compression schemes. Proceedings of the SPIE — The International Society for Optical Engineering, 2308:448–454, 1994. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Saupe:1994:BTC

[Sau94] Dietmar Saupe. Breaking the time complexity of fractal image compression. Technical Report 53, Institut für Informatik, Universität Freiburg, Freiburg, Germany, 1994. URL ftp://

                                        ftp.informatik.uni-freiburg.
                                        de/papers/fractal/Saup94a.
                                        ps.Z.

Saupe:1995:AFI

[Sau95a] Dietmar Saupe. Accelerating fractal image compression by multi-dimensional nearest neighbor search. In Storer and Cohn [SC95], pages 222–231. ISBN 0-8186-7012-6. ISSN 1068-0314. LCCN QA76.9.D33 D37 1995. URL ftp://ftp.

                                        informatik.uni-freiburg.de/
                                        papers/fractal/Saup95a.ps.

gz. to appear.

Saupe:1995:FIC

[Sau95b] Dietmar Saupe. Fractal image compression via nearest neighbor search. In Anonymous [Ano95], page ???? ISBN ???? LCCN ????

Saupe:1996:FSI

[Sau96a] Dietmar Saupe. The futility of square isometries in fractal image compression. In Proceedings ICIP-96 (IEEE International Conference on Image Processing), volume I, pages 161– 164. ????, ????, September 1996.

Saupe:1996:LDP

[Sau96b] Dietmar Saupe. Lean domain pools for fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2669:150–157, January 1996. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Storer:1993:DTD

[SC93] James A. Storer and Martin Cohn, editors. DCC’93: Third Data compression conference, March 30–April 2, 1993, Snowbird, Utah, USA. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3392-1, 0-8186-3391-3. ISSN 1068-0314. LCCN QA76.9.D33 D38 1993. IEEE catalog number 93TH0536-3. IEEE Computer Society Press order number 3392-02.

Storer:1995:DFD

[SC95] James A. Storer and Martin Cohn, editors. DCC’95: Fifth Data compression conference, March 30–April 2, 1993, Snowbird, Utah, USA. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7012-6. ISSN 1068-0314. LCCN QA76.9.D33 D37

Saupe:1994:CRM

[SH94] Dietmar Saupe and Raouf Hamzaoui. Complexity reduction methods for fractal image compression. In J. M. Blackledge, editor, Proceedings of the IMA Conference on Image Processing; Mathematical Methods and Applications. ????, ????, September 1994. URL ftp://ftp.

                                        informatik.uni-freiburg.de/
                                        papers/fractal/SaHa95a.ps.
                                        gz.

Saupe:1996:LAF

[SH96] Dietmar Saupe and Hannes Hartenstein. Lossless acceleration of fractal image compression by fast convolution. In Proceedings ICIP-96 (IEEE International Conference on Image Processing), volume I, pages 185– 188. ????, ????, September 1996.

Shonkwiler:1989:IAC

[Sho89] R. Shonkwiler. An image algorithm for computing the hausdorff distance efficiently. Information Processing Letters, 30:87– 89, January 1989. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Signes:1995:GII

[Sig95] Julien Signes. Geometrical interpretation of IFS based image coding. In Anonymous [Ano95], page ???? ISBN ???? LCCN ????

Signes:1996:RCS

[Sig96] Julien Signes. Reducing the codebook size in fractal image compression by geometrical analysis. Proceedings of the SPIE — The International Society for Optical Engineering, 2727:1400– 1409, 1996. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Simon:1995:ELB

[Sim95a] B. Simon. Explicit link between local fractal transform and multiresolution transform. In Proceedings ICIP-95 (IEEE International Conference on Image Processing), volume I, pages 278– 281. ????, ????, October 1995.

Simpson:1995:IPM

[Sim95b] Robert G. Simpson. Interband prediction method for subband image coding. Proceedings of the SPIE — The International Society for Optical Engineering, 2418: 60–66, 1995. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Stoksik:1995:PSA

[SLN95] Marc A. Stoksik, R. G. Lane, and D. T. Nguyen. Practical synthesis of accurate fractal images. Graphical models and image processing: GMIP, 57(3):206–219, May 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic).

Sloan:1994:LBR

[Slo94] Alan D. Sloan. Low-bit-rate fractal image coding. Proceedings of the SPIE — The International Society for Optical Engineering, 2239:210–213, 1994. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Shy:1994:FIC

[SLTC94] Chewn-Jye Shy, Hong-Yuan M. Liao, Chen-Kuo Tsao, and Ming-Yang Chern. Fractal image coding system based on an adaptive side-coupling quadtree structure. Proceedings of the SPIE — The International Society for Optical Engineering, 2308:455–466, 1994. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Saupe:1996:EFI

[SR96] Dietmar Saupe and Matthias Ruhl. Evolutionary fractal image compression. In Proceedings ICIP-96 (IEEE International Conference on Image Processing), volume I, pages 129– 132. ????, ????, September 1996.

Thomas:1995:RBF

[TD95] Lester Thomas and Farzin Deravi. Region-based fractal image compression using heuristic search. IEEE Transactions on Image Processing, 4(6):832–838, June 1995. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Titkov:1996:SBF

[TTMB96] Bronislav Titkov, Anatoli Tikhotskij, Alexandr Myboroda, and Helmut Buley. Structure-based fractal image coding. Proceedings of the SPIE — The International Society for Optical Engineering, 2952:652–663, 1996. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Tang:1994:FBI

[TW94] Yonghong Tang and William G.

Wee. Fractal-based image compression: a fast algorithm using wavelet transform. Proceedings of the SPIE — The International Society for Optical Engineering, 2308:1674–1682, 1994. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Walle:1995:RFI

[vdW95] Axel van de Walle. Relating fractal image compression to transform methods. Master’s thesis, University of Waterloo, Waterloo, Canada, 1995.

Walle:1996:MFI

[vdW96] Axel van de Walle. Merging fractal image compression and wavelet transform methods. In Anonymous [Ano95], page ???? ISBN ???? LCCN ????

Vines:1993:NAM

[VI93] Greg Vines and Monson H. Hayes III. Nonlinear address maps in a one-dimensional fractal model. IEEE Transactions on Signal Processing, 41(4): 1721–1724, April 1993. CO-DEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic). URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/Vine93.ps.gz.

Vines:1993:SMI

[Vin93] Greg Vines. Signal Modelling With Iterated Function Systems. PhD thesis, Georgia Institute of Technology, Atlanta, Georgia, May 1993. URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/Vine93.ps.gz.

Vines:1995:OBI

[Vin95] Greg Vines. Orthogonal basis IFS. In Yuval Fisher, editor, Fractal Image Compression: Theory and Application, chapter 10, pages 199–214. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94211-4.

Vences:1994:FCS

[VR94] Lucia Vences and Isaac Rudomin. Fractal compression of single images and image sequences using genetic algorithms. Technical report, Institute of Technology, University of Monterrey, 1994. URL ftp://ftp.

informatik.uni-freiburg.de/
papers/fractal/VeRu94.ps.gz.

Wohlberg:1994:RFI

[WdJ94] B. E. Wohlberg and G. de Jager. On the reduction of fractal image compression encoding time. In 1994 IEEE South African Symposium on Communications and Signal Processing (COMSIG ’94), pages 158– 161. ????, ????, October 1994.

URL http://www.dip.ee.uct.
ac.za/~brendt/publications/

wohlberg94.ps.gz. IEEE Cat

No 94TH0665-0.

Wohlberg:1995:FID

[WdJ95] B. E. Wohlberg and G. de Jager.

Fast image domain fractal compression by DCT domain block matching. Electronics Letters, 31 (11):869–870, May 1995. CO-DEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

URL http://www.dip.ee.uct.
ac.za/~brendt/publications/
wohlberg95.ps.gz.

Wohlberg:1996:FCP

[WdJ96] B. E. Wohlberg and G. de Jager.

Fractal coding performance for first-order Gauss–Markov models. Electronics Letters, 32(5): 441–442, February 1996. CO-DEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

URL http://www.dip.ee.uct.
ac.za/~brendt/publications/
wohlberg96a.ps.gz.

Wong:1993:FBI

[WHK93] Kwo-Jyr Wong, Ching-Han Hsu, and C.-C. J. Kuo. Fractal-based image coding with polyphase decomposition. Proceedings of the SPIE — The International Society for Optical Engineering, 2094:1207–1218, 1993. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Woolley:1994:RDP

[WM94] Stuart J. Woolley and Donald M.

Monro. Rate-distortion performance of fractal transforms for image compression. Fractals, 2(3 (or 6??)):395–398, 1994. CODEN FRACEG. ISSN 0218-348X.

                                        URL http://dmsun2.bath.ac.
                                        uk:8080/people/papers/montreal.
                                        ps.gz.

Woolley:1995:OPH

[WM95] Stuart J. Woolley and Donald M. Monro. Optimum parameters for hybrid fractal image coding. In Proceedings ICASSP-95 (IEEE International Conference on Acoustics, Speech and Signal Processing), volume 4, pages 2571–2574. ????, ????, May 1995.

Wilson:1994:RBF

[WNM94] D. L. Wilson, J. A. Nicholls, and D. M. Monro. Rate buffered fractal video. In Proceedings ICASSP-94 (IEEE International Conference on Acoustics, Speech and Signal Processing), volume 5, pages 505– 508. ????, ????, April 1994.

                                        URL http://dmsun2.bath.ac.
                                        uk:8080/people/papers/vid94.
                                        ps.gz.

Wohlberg:1996:FIC

[Woh96] Brendt Wohlberg. Fractal Image Compression and the Self-Affinity Assumption: A Stochastic Signal Modelling Perspective. PhD thesis, University of Cape Town, Cape Town, South Africa, August 1996. URL ftp:

                                        //ftp.dip.ee.uct.ac.za/pub/
                                        brendt/thesis.ps.gz.

Wood:1994:GDC

[Woo94] Christopher M. Wood. General data compression algorithm for space images using fractal techniques. Proceedings of the SPIE — The International Society for Optical Engineering, 2198:1336– 1341, 1994. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Wang:1996:PIF

[WY96] Zhou Wang and Ying-Lin Yu.

Partial iterated function systembased fractal image coding. Proceedings of the SPIE — The International Society for Optical Engineering, 2751:42–49, 1996. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Zhu:1994:FMO

[ZCT94] X. Zhu, B. Cheng, and D. M. Titterington. Fractal model of a onedimensional discrete signal and its implementation. IEE proceedings. Vision, image, and signal processing, 141(5):318–324, October 1994. CODEN IVIPEK. ISSN 1350-245x (print), 1359-7108 (electronic).

Zorpette:1988:FJA

[Zor88] Glenn Zorpette. Fractals: Not just another pretty picture. IEEE Spectrum, 25(10):29–31, October 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zhang:1995:FCI

[ZP95] Ying Zhang and Lai-Man Po.

Fractal color image compression using vector distortion measure. In Proceedings ICIP-95 (IEEE International Conference on Image Processing), volume III, pages 284–287. IEEE Signal Processing Society, Washington, DC, USA, October 1995.

Zhang:1991:HIC

[ZY91] N. Zhang and H. Yan. Hybrid image compression method based on fractal geometry. Electronics Letters, 27(5):406–408, February 1991. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Zhao:1994:ICU

[ZY94] Y. Zhao and B. Yuan. Image compression using fractals and discrete cosine transform. Electronics Letters, 30(6):474–475, March 1994. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Zhang:1996:HFI

[ZZZ^{+}96] Zhengbing Zhang, Yaoting Zhu, Guangxi Zhu, Hanqiang Cao, and Donghui Xue. Hybrid fractal image coding method. Proceedings of the SPIE — The International Society for Optical Engineering, 2727:1360–1366, 1996. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).